Based upon how the titration of the amino acid was performed, and the data that you collected from the titration, calculate the concentration of the amino acid that was titrated. NOTE: you cannot base your calculation upon the assumption that the amino acids were fully protonated at the start of the titration. When you calculate the volume of titrant used you cannot start at 0.0mL. This is because the amino acids solutions are prepared by the addition of concentrated HCl - a step which is subject to measurement error. This means that the amino acid may not be fully protonated at the start of the titration if too little HCl has been added, or if too much HCl was added, those protons need to be neutralized before you begin titrating off protons from the amino acid.

Titrated 30.0mL of unknown amino acid with 150mM NaOH.

pKa for amino acids are: 2.3 (carboxyl) and 9.7 (amino)
Answer #1

In the presence of HCl, if there is any COO- moiety present in amino acid, it will become COOH, leaving the NH3+ moiety as such.

i.e. There are two acidic protons in the amino acid, which correspond to COOH as well as NH3+.

Therefore, you need two moles of a base to titrate one mole of amino acid.

The question should give the volume of 150 mM NaOH solution!

For example, if the volume of 150 mM NaOH solution is 10 mL, then the millimoles of NaOH = 150 mmol/L * (10/1000) L = 1.5 mmol

For the complete titration, the millimoles of amino acid = 1.5/2 = 0.75 mmol

Now, the concentration of amino acid = 0.75 mmol/(30/1000) L = 25 mM